ΔHf/ KJ mol-1:
Mg(s) and O2(g) 0
MgO(s) - 601.70

Cp, m / KJ mol-1
Mg - 24.89
O2 - 29.36
MgO - 37.15

Define the std enthalpy of combustion and calculate this parameter for the combustion of magnesium at 25 degrees celsius.

Provide a balanced chemical equation, the equation used to calculate ΔHc and finally the complete thermochemical equation for the reaction.
1
Expert's answer
2019-07-09T02:00:42-0400

From van't hoff's law:

Standard enthalpy is difference between products and substrats.

balanced chemical equation:

2Mg+O2 = 2MgO

ΔHc = ΔHp - ΔHs = -601.7 - 0 = --601.7 kJ/mol


complete thermochemical equation :

2Mg+O2 = 2MgO +1203.4 kJ